OpenStudy (anonymous):

Give an example of three numbers the fit the description below: - the sum of the numbers is positive and the product is negative

OpenStudy (anonymous):

Ok so with 3 numbers there are 2 ways to get a negative product. 1. - * - * - = - Or 2. + * + * - = - Since we know the sum is going to be positive, we are going to use the 2nd example. So we need to find 3 numbers that the 2 positives together are greater than the negative.

OpenStudy (anonymous):

So, just for an extreme example, 1000 + 1000 + (-1) are 3 numbers that have a positive sum. (1999)

OpenStudy (anonymous):

But have a product, 1000 * 1000 * (-1) = -1,000,000

OpenStudy (anonymous):

That is negative. So that being said, can you now find two more examples that multiply into negative but add into positive?

OpenStudy (dan815):

-1+2+3=4 product -1*2*3=-6 think of more

OpenStudy (anonymous):

Thanks a lot guys, I appreciate it! :)